umerlodhi (umerlodhi):

9th grade math pls help

OpenStudy (4everaddicted2anime):

What is the question?

umerlodhi (umerlodhi):

the answers

OpenStudy (welshfella):

y = mx + b m = slope and b = intersect on the y axis

OpenStudy (4everaddicted2anime):

Using the coordinates (0,1) and (3,5), can you find the slope?

umerlodhi (umerlodhi):

no

OpenStudy (4everaddicted2anime):

slope formula is \[\frac{ y _{2}-y_{1} }{ x_{2}-x_{1} }\]

umerlodhi (umerlodhi):

can you please just say the answer

OpenStudy (4everaddicted2anime):

\[y=\frac{ 4 }{ 3 }x+1\]

OpenStudy (anonymous):

y=4/3x+1

OpenStudy (anonymous):

geometry?

umerlodhi (umerlodhi):

thanks maybe this can help u https://mathway.com/ yes geometry

OpenStudy (anonymous):

if it is, don't ask me a proving question... i hate em :p

umerlodhi (umerlodhi):

okay i have a few more